OpenStudy (ksaimouli):

y'+y=0 find a general solution for the homogeneous differential equations

OpenStudy (ksaimouli):

^ is that linear?

OpenStudy (vincent-lyon.fr):

Yes it is, because if y1 and y2 are solutions, then \(\lambda\, y_1+\mu\, y_2\) is a solution too.

OpenStudy (ksaimouli):

It is linear so I used integration factor

OpenStudy (ksaimouli):

\[\frac{ dy }{ dt }+a(t)y=0\]

OpenStudy (ksaimouli):

\[e^{A(t)}; A(t)=\int\limits_{}^{}1 dt\]

OpenStudy (ksaimouli):

so I got \[Ce^t\]

OpenStudy (ksaimouli):

but the answer is \[Ce^{-t}\]

OpenStudy (ksaimouli):

@phi

OpenStudy (phi):

\[ \frac{ dy }{ dt }+y =0\] multiply by the int. factor: \[ e^t\frac{ dy }{ dt }+e^ty=0 \\ \frac{ d }{ dt }\left( e^ty\right) = 0 \\ e^ty = C\\ y = Ce^{-t} \]

OpenStudy (ksaimouli):

got you

OpenStudy (ksaimouli):

@phi shouldn't the C be -C or it just don't matter?

OpenStudy (phi):

C is an unknown constant. It could be plus or minus, but unless we are given more info, we don't know. -C is ok, but unless you have a good reason to toss in extra symbols, I would leave it just C)